OpenStudy (anonymous):

what is 2(p+3)-8p+13

OpenStudy (zkoqmosii):

Do you want it simplified?

OpenStudy (anonymous):

yes i forgot to add that thanks

OpenStudy (zkoqmosii):

2 (p+3)-8 p+13 Distribute 2 over p+3. 2 (p+3) = 2 p+6: 2 p+6-8 p+13 Group like terms in 2 p-8 p+6+13. Grouping like terms, 2 p-8 p+6+13 = (6+13)+(2 p-8 p): (6+13)+(2 p-8 p) Combine like terms in 2 p-8 p. 2 p-8 p = -6 p: -6 p+(6+13) Evaluate 6+13. 6+13 = 19: Answer: 19-6 p

OpenStudy (anonymous):

wow thank you